### v1

# Surajit Das The code will be provided upon request to the corresponding author.11 0009-0008-6692-6697    Pavel Zun 22 1111-2222-3333-$44$

Abstract
This article presents a novel microscopy image analysis framework designed for low-budget labs equipped with a standard CPU desktop. The Python-based program enables cytometric analysis of live, unstained cells in culture through an advanced computer vision and machine learning pipeline. Crucially, the framework operates on label-free data, requiring no manually annotated training data or training phase. It is accessible via a user-friendly, cross-platform GUI that requires no programming skills, while also providing a scripting interface for programmatic control and integration by developers. The end-to-end workflow performs semantic and instance segmentation, feature extraction, analysis, evaluation, and automated report generation. Its modular architecture supports easy maintenance and flexible integration while supporting both single-image and batch processing. Validated on several unstained cell types from the public dataset of livecells, the framework demonstrates superior accuracy and reproducibility compared to contemporary tools like Cellpose and StarDist. Its competitive segmentation speed on a CPU-based platform highlights its significant potential for basic research and clinical applications—particularly in cell transplantation for personalized medicine and muscle regeneration therapies.
Keywords:
Segmentation Graphical User Interface Cell Culture Machine learning Computer vision
1 Introduction
This article outlines the core features [^1] The algorithmic discussion is incorporated in separated article. of an integrated Python-based image analysis GUI application for biomedical research or clinical practice, focusing on cellular and morphological characterization. It provides image inspection, processing, and quantitative analysis through a graphical user interface (GUI).
Image Management and Visualization: The tool supports standard formats (PNG, JPG, BMP, TIFF) with an interactive canvas for scrolling, zooming, and background opacity adjustment for overlays. Users can define regions-of-interest (ROI) for independent analysis.
Interactive Pixel and Spatial Analysis: Real-time cursor tracking displays XY coordinates. Pixel inspection retrieves RGB values and grayscale conversions. User-defined neighborhood kernels enable first-order (mean, median, mode, SD), higher-order (skewness, kurtosis), and spatial statistics (Moran’s I, variogram measures). Results appear dynamically in a scrollable widget.
Modular Image Processing Pipeline: The GUI links to a modular workflow focused on segmentation: (I) Semantic Segmentation: Primary segmentation (‘HIP Actual Grayscale Analysis’) with configurable parameters (shift, span, masking threshold, contrast threshold, randomness index). (II) Post- Segmentation: Refinement via denoising/cleaning with morphological operations. (III) Instance Segmentation: Algorithms for separating individual objects (e.g., cells). (IV) Analysis & Reports: Extraction of features and auto-generation of reports.
Comprehensive Quantitative Analysis: Users can compute a wide range of metrics: General (object count, cell density); Geometry (perimeter, area, aspect ratio); Morphology (roundness, shape, convexity); Localization (2D coordinates, centroid); Orientation (object orientation); and Voronoi Entropy (spatial disorder from Voronoi tessellation).
Data Handling and Reporting: Pixel statistics and analysis results export to Excel. The ‘Report’ function generates analytical profiles with a progress bar for long tasks.
Profile-Based Batch Processing: Profiles store reusable parameters for single or batch processing, ensuring reproducibility. Batch mode produces analytical reports with statistical summaries and visualizations, suitable for research and clinical use.
2 Literature Review
Over the past two decades, microscopy image analysis has progressed from manual thresholding to AI-driven methods, enhancing accuracy, reproducibility, and throughput. Early tools like CellProfiler (2006) [1], ImageJ/Fiji [2], and Icy (2011) [3] combined classical image processing with basic ML for segmentation and tracking but relied on staining, high-contrast images, and manual tuning.
The 2010s saw ML-based tools such as Ilastik (2011) [4], CellCognition (2010) [5], and TrackMate (2017) [6], enabling pixel classification, supervised learning, and improved tracking. Feature engineering was labor-intensive, and performance varied with image quality.
Deep learning (2018–present) transformed segmentation. DeepCell (2016) [7], Cellpose (2020) [8], and StarDist (2018) [9] achieved near-human accuracy. Cellpose, using a modified U-Net, predicts horizontal and vertical flow fields along with binary cell masks, which are post-processed for precise instance segmentation, even for irregularly shaped cells. ZeroCostDL4Mic (2020) [12] enabled DL without coding, though requiring large datasets and high computation.
Recent tools like Omnipose (2022) [10] and MISIC (2023) [11] handle complex shapes, unstained cells, and sparse label conditions. Omnipose excels in bacterial and irregular cells; MISIC in stain-free and sparse-label scenarios. Emerging trends include hybrid classical-DL models, self-supervised learning, and user-friendly GUIs. Remaining gaps include domain-specific analysis (e.g., myoblasts), multi-modal integration, computational cost, and accessibility. The proposed framework addresses these gaps via classical processing, spatial statistics, and DL in a Python-based, user-friendly tool for regenerative medicine.
3 In-Depth Functionalities
3.1 Segmentation
The semantic segmentation module applies neighborhood-based statistical filtering and fuzzy intensity transformation to separate texture (cells) from background. It integrates Numba-optimized functions for efficiency, using local statistics (mean, median, mode, skewness, kurtosis), spatial measures (Moran’s I, variogram, adjacency shift), and fuzzy logic for pixel classification. Additional steps include homogeneous masking, circle removal via Hough transform, and partially adaptive thresholds (sv1–sv5) etc. Processing is performed with multi-channel support, producing binary masks that serve as inputs for post-processing and instance analysis. Users can execute the segmentation task either through the GUI (via a click) or via the console commands.
Figure 1: GUI for Sampling Background Pixels with neighbourhood $5\times 5$
Figure 2: (Left) Calibration Window, (Right) Denoising Window.
3.2 Post-Segmentation Processing Module
The module provides a comprehensive suite for refining and denoising segmented images.
Core Architecture: Implemented as a ProfileProcessor class managing the post-processing pipeline. Key data structures include: image data (BGR), contour list, operations history, and current profile. SQLite ensures persistent storage of operation history with timestamps, parameters, and performance metrics for auditability.
Input/Output Handling: Supports loading pickle files with images, contours, histories, and profiles. Output is saved as both pickle (for further processing) and PNG (for visualization), with proper color space conversion between grayscale and BGR.
Contour Processing and Analysis: Enables contour detection and management with intelligent sampling (configurable spacing, Euclidean-distance-based point reduction, structured sampling) while preserving morphological integrity. Fully compatible with OpenCV contour functions.
Morphological Operations: - Area-based Filtering: Eliminates regions below minimum area, with modes for removal (filling) or marking (visualization); supports hierarchical contour processing. - Shape-based Filtering: Aspect ratio elimination (customizable bounds), circularity filtering (isoperimetric quotient), eccentricity-based filtering for elliptical structures. - Image Enhancement: Median blur (auto kernel size), erosion/dilation (configurable kernels), and threshold-based filling for artifact removal.
Advanced Contour Refinement: Includes Sobel gradient edge detection, distance transforms, force-based contour deformation (parameters: force constant, decay exponent, gradient threshold), convex hull enforcement, and multi-stage smoothing via neighborhood averaging.
Parameter Normalization System: Provides explicit mapping of parameter names, automatic key normalization, numeric type conversion, and flexible matching strategies for consistent operation handling.
Operation Mapping and Execution: Supports multiple naming conventions, case-insensitive and substring-based matching, parameter validation, adaptive execution, and robust error handling with fallbacks.
Quality Control and Metrics: Features exception handling with detailed reporting, success/failure tracking, performance metrics recording, and full processing history maintenance. This module offers a complete solution for post-segmentation refinement and analysis of biomedical images, particularly for cellular morphology studies requiring precise contour characterization and filtering for accurate quantitative analysis. Users can execute "Post-Segmentation Processing" either through the GUI (via a click) or via the console commands.
3.3 Instance Segmentation and Morphological Analysis Module
This module automates detection, segmentation, and quantitative analysis of individual objects, focusing on cellular structures.
Core Functionality: Performs instance segmentation via contour-based detection, extracting features and generating annotated visuals and quantitative data.
Input Processing: Accepts images in multiple formats via pickle serialization (grayscale, single-channel BGR, three-channel BGR). Automatic color conversion ensures consistency; adaptive thresholding generates binary masks for contour detection.
Contour Detection and Object Identification: Uses OpenCV contour detection with , configurable minimum area thresholds remove artifacts, and each object receives a unique identifier.
Morphological Feature Extraction: Computes geometric features (area, perimeter, aspect ratio, bounding box), morphological metrics (roundness, shape classification, convexity, solidity), spatial/orientation features (centroid, top-left coordinates, orientation angle, eccentricity), and advanced metrics (Voronoi entropy, cell density, equivalent diameter, extent).
Visual Output Generation: Produces annotated, bounding-box, and contour images with configurable annotation parameters (font scale, color, thickness).
Data Export and Reporting: Exports structured Excel spreadsheets with all features, automatic naming, console summaries, real-time feedback, and debugging support.
Configuration and Customization: Users can adjust feature selection, minimum area thresholds, annotation appearance, and output management (file naming, directories).
Scientific Image Analysis: Enhances extended morphological characterization supporting intensity-based features, advanced convex hull metrics, refined eccentricity and solidity, and principal axis analysis for orientation.
Performance and Reliability: Robust error handling, input validation, adaptive processing, memory-efficient contour handling, and cross-platform compatibility. This module provides a reliable solution for automated cellular analysis, combining quantitative data with visual verification for biomedical morphological studies.
Figure 3: (Left) Profile selection, (Right) Report
3.4 Automated Report Generation Module
This module produces comprehensive HTML reports from Excel-based cellular analysis data, completing the image analysis pipeline by converting quantitative morphological measurements into structured, presentable outputs.
Core Functionality: Automatically generates HTML reports with statistical summaries, distribution visualizations, structured analytical results, and professional formatting for research documentation.
Architecture: Built on jitprofiler with modular separation of data processing and presentation layers.
Inputs: Excel files containing geometric, morphological, and spatial measurements, along with configuration parameters (image identifiers and processing settings).
Outputs: Professional HTML documents with structured layouts, integrated charts, descriptive statistics, and cross-platform compatibility.
File Management: Standardized naming conventions and directories embed image identifiers, neighborhood size, timestamps, or version indicators for traceability.
Pipeline Integration: Consumes processed data from segmentation and feature extraction, maintaining parameter consistency and ensuring quality through standardized reporting.
Error Handling: Validates file existence and format, checks parameter integrity, and provides informative messages. where \<input\_image\_path> is the source image and \<neighborhood\_size> the feature extraction parameter.
Quality Assurance: Ensures input completeness, parameter-result consistency, standardized formatting, and cross-referencing for multi-image studies.
The module transforms raw cellular data into actionable insights and polished research documentation. Users can execute Report Generation task either through the GUI (via a click) or via the console commands.
3.5 Profile Management and Batch Processing System
The profile manager allows researchers to save, reuse, and share complete image processing configurations, ensuring reproducibility and consistency across experiments. Profiles are stored in a standardized JSON format with metadata (creation time, source image, version) and all analysis parameters. A graphical interface provides an interactive browser for managing profiles, previewing settings, and configuring directories, while the batch engine automates large-scale processing with real-time progress updates, error handling, and cancellation options.
From the user’s perspective, workflows are straightforward: create a profile once, apply it to single images or entire datasets, and receive consistent results with automated reporting. The system manages input/output directories automatically, organizes results into structured folders, and preserves original naming for easy traceability.
Key benefits include reproducibility (version-controlled profiles and audit trails), scalability (batch execution on large collections or time-series data), and flexibility (parameter optimization, protocol standardization, or quality control). Built-in error recovery, efficient memory handling, and detailed logging ensure reliability even in high-throughput studies.
In practice, the system streamlines experimental pipelines: users define their analysis settings once, let the system run across thousands of images, and obtain consistent, validated outputs with minimal manual intervention.
4 Results:
Fig 4 displays the output against a sample image taken from public dataset of livecell [^2] https://sartorius-research.github.io/LIVECell/ (image name: A172_Phase_A7_1_00d00h00m_3.PNG). The semantic segmentation image achieved Accuracy: 0.9759, Precision: 0.8613, Recall: 0.9482, F1: 0.9027, Dice: 0.9027, Jaccard: 0.8226.
However, for different categories the scores differs. We mention the average scores of 3178 images. Complete results are available at 333 https://drive.google.com/file/d/1UWfLJ1Zg263zppYRzza3OC9rI14q2qmO/view?usp=sharing.
Table 1: Dataset of 3178 images: Segmentation performance metrics
Dice IoU Accuracy Precision Recall F1 Score SSIM Hausdorff
Mean 0.814431 0.694019 0.872015 0.803413 0.851245 0.814431 0.502930 57.287395
Std 0.078670 0.106993 0.096104 0.112562 0.124639 0.078670 0.245851 34.920233
(a) Input Image
(b) Annotated Image
(c) Bounding Box
(d) Detected Contours
Figure 4: Image processing results
Conclusion
In summary, this software tool provides a unified platform for the entire image analysis workflow, from initial image inspection and interactive data exploration to advanced, configurable segmentation and quantitative morphological analysis. Its design is particularly suited for biomedical research applications requiring detailed characterization of cellular structures and spatial relationships.
{credits}
4.0.1 Acknowledgements
The research was supported by ITMO University Research Projects in AI Initiative (RPAII) (project #640103, Development of methods for automated processing and analysis of optical and atomic force microscopy images using machine learning techniques)*.
[…]
The authors declare that they have no competing interests relevant to the content of this article.acknowledgments [^4] This preprint is not peer-reviewed and is provided under a specific license. Unauthorized use is prohibited and is the sole responsibility of the user.,